OpenStudy (anonymous):

for the pair of fnctions, f(x) = x+2 and g(x)=x^2+x-7, find the following a) (f.g)(x) b) (f.g)(3) c) (g.f)(x) d) (g.f)(3)

OpenStudy (anonymous):

I am correct that f.g is multiplication?

OpenStudy (anonymous):

a)\[(f*g)(x)=f(x)*g(x)=(x+2)(x^2+x-7)=x^3+3x^2-5x-14\] b)\[(f*g)(3)=3^3+3*3^2-5(3)-14=25\] c) \[(g*f)(x)=(f*g)(x)\] d)\[(g*f)(3)=(f*g)(3)=25\]